### v2

# A Comprehensive Analysis of Social Tie Strength: Definitions, Prediction Methods, and Future Directions

Abstract
The rapid growth of online social networks has underscored the importance of understanding the intensity of user relationships, referred to as “tie strength.” Over the past few decades, extensive efforts have been made to assess tie strength in networks. However, the lack of ground-truth tie strength labels and the differing perspectives on tie strength among researchers have complicated the development of effective prediction methods for real-world applications. In our study, we first categorize mainstream understandings of tie strength into seven standardized definitions and verify their effectiveness by investigating the class distributions and correlations across these definitions. We also draw key insights into tie resilience from the perspective of tie dissolution that (1) stronger ties are more resilient than weaker ones, and (2) this tie resiliency ratio increases as the network evolves. We then conduct extensive experiments to evaluate existing tie strength prediction methods under these definitions, revealing that (1) neural network methods capable of learning from semantic features hold great potential for high performance, (2) models struggle under definitions that offer limited understandings of tie strength in the network, (3) existing models face imbalance issues that cannot be addressed by traditional quantity imbalance techniques, and (4) different definitions of tie strength allow for the inference of not only the current state but also the future state of a tie. Building on these findings, we propose strategies to improve existing methods and suggest several promising directions for future research.
1 Introduction
Recent years have witnessed the explosive growth of social networks. One of the most important components of social networks is the relationship between users, and the intensity of these relationships is often referred to as “tie strength” (Marsden and Campbell 1984). Tie strength has been widely studied and can have various definitions depending on the context and interpretation. For example, the tie strength between two users on Twitter (X) may be considered high if there is a bidirectional relation between them as it would indicate mutual interest. In another definition, two users may have a high tie strength if they frequently “like” or interact with each other’s posts on the platform. Extensive research has demonstrated the importance of understanding tie strength within social networks as it can extract meaningful insights from network data (Sintos and Tsaparas 2014). For instance, tie strength is a crucial factor in link prediction within networks (Li et al. 2013; Yang et al. 2019). Stronger ties in business networks are often correlated with successful trades, leading to better financial understanding (Kang et al. 2021; Wong et al. 2022). Additionally, tie strength helps predict epidemic and pandemic spread, aiding in disease prevention and improving healthcare planning (Krendl and Perry 2021; Birditt et al. 2021).
During the past few decades, methods to infer/predict tie strength in social networks have been studied. The initial research (Granovetter 1973; Marsden and Campbell 1984; Petróczi, Nepusz, and Bazsó 2007; Gilbert and Karahalios 2009) developed heuristic methods to measure tie strength based on feedback from interviewees. Later researchers began to investigate using non-structural indicators such as frequency and duration of interaction (Zhuang et al. 2021; Jones et al. 2013; Xiang, Neville, and Rogati 2010), level of interaction or emotional intensity and intimacy (Gilbert and Karahalios 2009; Liu, Li, and Wang 2020), and the reciprocal services found within the tie (Sintos and Tsaparas 2014; Adriaens et al. 2018; Zhong et al. 2020; Oettershagen, Konstantinidis, and Italiano 2022). However, these approaches generally rely on social network user characteristics that are often incomplete and difficult to measure (Adriaens et al. 2018). To address this issue, several methods have been developed that either rely solely on the Strong Triadic Closure (STC) principle (Sintos and Tsaparas 2014; Adriaens et al. 2018; Oettershagen, Konstantinidis, and Italiano 2022) or use heuristic approaches based on topological properties such as node degrees (Kahanda and Neville 2009), clustering coefficients (Mattie et al. 2018), or the overlap between nodes’ neighborhoods (Onnela et al. 2007) to characterize tie strength. In recent years, with the emergence of network embedding techniques, methods that generate embeddings to reflect the latent network structures properties have also been proposed (Zhong et al. 2020; Zhuang et al. 2021).
However, ground-truth tie strength label information is usually missing or incomplete in practice. Most online social network platforms lack mechanisms to distinguish between different types of ties. Furthermore, obtaining ground-truth labels of tie strengths in networks is an under-explored problem (Liu, Li, and Wang 2020). Existing literature has various tie strength definitions, including direction-based definitions (Sintos and Tsaparas 2014; Adriaens et al. 2018; Zhong et al. 2020; Oettershagen, Konstantinidis, and Italiano 2022), score-based definitions (Liu, Li, and Wang 2020) and frequency-based definitions (Zhuang et al. 2021). Since the differences between these definitions vary greatly, the tie strength prediction performance of these methods under different definitions also differs significantly. Therefore, it is necessary to devise a uniform definition of ground-truth tie strength labels and develop corresponding methods. As the field evolves, consistency across studies will enable more meaningful comparisons and allow future research to build on existing work more effectively.
As suggested by the empirical results later Section 5, existing models tend to have low prediction performance and insignificant average $tie$ weight differences between strong and weak ties. This can significantly affect the models’ ability to generalize across various real-world scenarios, and limit their effectiveness in diverse settings. In addition, failure to correctly recognize strong and weak ties can have significant implications in real-world networks, such as underestimating the influence of fraudulent behavior, or miscalculating the risk of a community suffering from a disease. It is important to understand the reasons why current tie strength prediction models fail to perform well under various tie strength definitions, and consider how we can obtain better predictions under weakly-/un-supervised settings.
To tackle these challenges, we first summarize the mainstream understandings of tie strength into seven standardized definitions and validate their significance by examining class distributions and correlations. Additionally, we draw valuable insights into tie resilience from the perspective of tie dissolution: (1) strong ties are more resilient than weak ties, and (2) this tie resiliency ratio increases as the network evolves. We then carried out extensive experiments to evaluate the performance of existing methods, leading to several key findings: (1) neural network models that effectively use semantic features have strong potential for high performance, (2) models face difficulties when using definitions that offer limited insights into tie strength, (3) current methods struggle with imbalance issues that cannot be addressed by traditional techniques, and (4) different definitions of tie strength allow for the prediction of different states of that tie in a dynamic network. Based on these findings, we propose strategies to improve existing models and suggest several promising directions for future work in tie strength prediction. The contributions of this study are as follows:
  • We summarize the mainstream practices for assigning tie strength labels and develop 7 standardized definitions, which we use as pseudo ground-truth tie strength labels.
  • We analyze the tie strength definitions, including correlation analysis and resilience of strong/weak ties from the perspective of tie dissolution.
  • We conduct experiments on tie strength prediction methods across numerous settings and develop new understandings from various perspectives.
  • We propose improvement strategies based on our experiments to highlight multiple future research directions.
The paper is structured as follows: We review existing tie strength prediction works and formalize the problem settings in Section 2. We revisit mainstream understandings of tie strength, and propose and investigate seven standardized definitions, including from a tie resilience perspective in Section 3. In Section 4, we present real-world settings and prediction methods. We provide extensive experiments, analyzing model performance across definitions and proposing strategies for improvement in Section 5. Finally, Section 7 concludes the paper.
2 Preliminaries
In this section, we first provide an overview of prior work focusing on measuring and predicting tie strength. Then we introduce notations and tie strength prediction settings.
2.1 Related Work
With the development of online social networks, tie strength prediction has received increasing attention  (Marsden and Campbell 2012; Gilbert 2012). While recent papers (Adriaens et al. 2018; Mattie et al. 2018) have explored tie strength at fine-grained levels, the most fundamental and widely accepted understanding (Granovetter 1973; Gilbert and Karahalios 2009; Kumar, Rao, and Nagpal 2012) favors the binary categorization of ties into strong/weak categories.
However, there is no standard way to obtain the ground truth tie strength (Liu, Li, and Wang 2020), and most online platforms even lack mechanisms to distinguish between different types of ties. To fill this gap, some of the pioneering work (Petróczi, Nepusz, and Bazsó 2007) relies on the feedback from questionnaires based on the theory proposed by  (Granovetter 1973), that is tie strength is a combination of the amount of time, the emotional intensity, the intimacy (mutual confiding) and reciprocal services.
Some other works also use semantic features to infer tie strength. For example,  (Xiang, Neville, and Rogati 2010) proposed a model to infer relationship strength based on profile similarity and interaction activity. However, such approaches rely heavily on the feedback qualities of users and, hence are hard to use for large social networks. To solve the issues, recent researchers began to develop heuristic methods based on structural features. For example,  (Mattie et al. 2018) developed heuristic methods that used structure features such as degree, clustering coefficients, unweighted overlap, etc as predictors.  (Sintos and Tsaparas 2014; Adriaens et al. 2018; Zhong et al. 2020; Oettershagen, Konstantinidis, and Italiano 2022) also use strong triad closure properties (Granovetter 1973) to infer tie strength. Additionally, with the development of neural networks,  (Liu, Li, and Wang 2020; Zhuang et al. 2021) developed network embedding techniques to infer tie strength.
Despite these efforts, existing research poses different perspectives towards strong/weak ties. Some researchers  (Sintos and Tsaparas 2014; Adriaens et al. 2018; Zhong et al. 2020; Oettershagen, Konstantinidis, and Italiano 2022) built their methods based on directional-based definitions that assume bidirectional edges are strong ties, while other researchers (Zhuang et al. 2021; Liu, Li, and Wang 2020; Zhuang et al. 2021; Jones et al. 2013; Xiang, Neville, and Rogati 2010) use score or frequency criteria to determine the ground-truth tie strength. To the best of our knowledge, no previous research has investigated these understandings and how different models perform under various perspectives, which makes it challenging to ensure solid performance from existing models in different real-world scenarios. To deepen our understanding of tie strength and shed light on future directions, we explicitly conduct this paper to comprehensively study the performance of existing tie strength methods under different tie strength definitions.
2.2 Notations
Let $G=(\mathcal{V},\mathcal{E},\mathbf{X},\mathbf{E},\mathbf{W})$ be a directed attributed network, where $\mathcal{V}=\{v_{i}\}_{i=1}^{n}$ is the set of $n$ nodes (i.e., $n=|\mathcal{V}|$) and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of $m$ edges (i.e., $m=|\mathcal{E}|$) with $e_{ij}$ denoting the edge between the node $v_{i}$ and $v_{j}$. $\mathbf{X}\in\mathbb{R}^{n\times d}$ represents the node feature matrix with dimension $d$. $\mathbf{E}\in\mathbb{R}^{m\times d^{\prime}}$ represents the edge feature matrix with dimension $d^{\prime}$. The edge weight matrix is $\mathbf{W}\in\mathbb{R}^{n\times n}$ where $\mathbf{W}_{ij}$ is the weight of $e_{ij}\in\mathcal{E}$. The adjacency matrix is denoted as $\mathbf{A}\in\{0,1\}^{n\times n}$ with $\mathbf{A}_{ij}=1$ if an observed edge exists from node $v_{i}$ to $v_{j}$ and $\mathbf{A}_{ij}=0$ otherwise.
The diagonal matrix of node degree is notated as $\mathbf{D}\in\mathbb{Z}^{n\times n}$ with the degree of node $v_{i}$ being $d_{i}=\mathbf{D}_{ii}=\sum_{j=1}^{n}\mathbf{A}_{ij}$. The neighborhood node set of node $v_{i}$ is denoted as $\mathcal{N}_{i}$.
2.3 Tie Strength Prediction Problem Settings
In the tie strength prediction tasks, let the (pseudo) ground-truth tie strength label be $\mathbf{Y}\in\{0,1\}^{n\times n\times C}$, where $C$ is the total number of discrete tie strength categories, and $\{\mathcal{E}_{k}\}_{k=1}^{C}$ is the set of labeled ties with category $k$. Note that since tie strength is a reflection of the intensity of reciprocal relationships between users, a tie is undirected, which means there is only one tie strength between node $v_{i}$ and node $v_{j}$. Then:
  • Weakly-supervised settings: Given a directed network $G$ with training tie strength set $\mathbf{Y}^{\text{Tr}}$, the task aims to correctly infer the strength of ties between nodes in the $\mathbf{Y}^{\text{Test}}$.
  • Unsupervised settings: Given a directed network $G=$ with no training tie strength set, the task aims to correctly infer the strength of ties between nodes in the $\mathbf{Y}$.
Note that in this study we only consider binary tie strength (i.e., $C=2$) where existing edges can be either strong ($S$) or weak ($W$), which is a common setting in this line of research (Liu, Li, and Wang 2020).
3 Investigating Tie Strength Definitions
In this section, we start by analyzing the existing tie strength definitions and organize them into 3 main categories. Then we propose a set of 7 tie strength definitions that span the primary tie strength categories from existing literature. To develop a more comprehensive understanding of these definitions, we collect three real-world datasets to conduct various analyses. Specifically, this includes analyzing the strong/weak tie distributions, pairwise correlation of strong/weak labeling of edges between definitions, and the resilience of ties from the perspective of tie dissolution.
3.1 Tie Strength Definitions
The early work related to tie strength often relies on feedback from users who agree to participate in a survey (Petróczi, Nepusz, and Bazsó 2007; Gilbert and Karahalios 2009). However, this approach heavily relies on the feedback quality and is hard to generalize to large-scale social networks. In recent years, researchers have started to develop different methods to assign labels of tie strength based on different application scenarios. Most of the efforts in this direction can be roughly categorized into three categories:
  • Direction-based Definitions: As suggested in  (Sintos and Tsaparas 2014; Adriaens et al. 2018; Zhong et al. 2020; Oettershagen, Konstantinidis, and Italiano 2022), a “strong tie” can be understood as a bidirectional/reciprocal relationship between users, as it would imply that the users know each other.
  • Score-based Definitions: Another approach is to assign tie strength labels based on the score between users. For example,  (Liu, Li, and Wang 2020) proposed assigning tie strength based on rating scores (which represent the level of trust between buyers/sellers) exchanged between anonymous users in the Bitcoin networks.
  • Frequency-based Definitions: Many works also assign tie strength labels based on frequency-related information. For instance,  (Zhuang et al. 2021; Jones et al. 2013; Xiang, Neville, and Rogati 2010) all consider tie strength through the lens of how frequently users interact.
In most real-world datasets, both score-related and frequency-related information can usually be reflected in edge weight (i.e., $\mathbf{W}$). For example, the edge weight in Bitcoin_alpha (Derr et al. 2017) datasets represents the trust level rating score from one user to another, and the edge weight in the CollegeMSG (Panzarasa, Opsahl, and Carley 2009) dataset represents the contact frequency from one user to another. Hence, we summarized the existing understandings of tie strength into 7 definitions based on the edge weight and edge direction between users. More specifically, strong edges (i.e., $\mathbf{Y}_{ijS}=1$) are assigned according to the following conditions:
  • I: if bidirectional between $v_{i}$ and $v_{j}$, i.e, $e_{ij},e_{ji}\in\mathcal{E}$.
  • II: if $\mathbf{W}_{ij}+\mathbf{W}_{ji}\geq\Theta$.
  • III: if bidirectional and $\mathbf{W}_{ij}+\mathbf{W}_{ji}\geq\Theta$.
  • IV: if both $\mathbf{W}_{ij}\geq\mathbf{W}_{i}(\theta)$ and $\mathbf{W}_{ji}\geq\mathbf{W}_{j}(\theta)$.
  • V: if either $\mathbf{W}_{ij}\geq\mathbf{W}_{i}(\theta)$ or $\mathbf{W}_{ji}\geq\mathbf{W}_{j}(\theta)$.
  • VI: if $\mathbf{W}_{ij}+\mathbf{W}_{ji}\geq\mathbf{W}_{i}(\theta)+\mathbf{W}_{j}(\theta)$.
  • VII: if bidirectional and $\mathbf{W}_{ij}+\mathbf{W}_{ji}\geq\mathbf{W}_{i}(\theta)+\mathbf{W}_{j}(\theta)$.
Table 1: Comparison among definitions of strong ties based on their bidirectionality and weight information.
Tie Strength Definitions
I II III IV V VI VII
Bidirectional
Global Weight Sum Sum
Local Weight And Or Sum Sum
In the above definitions, $\Theta\in(0,\infty)$ is the global threshold, and $\mathbf{W}_{i}(\theta)$ is the top $\theta$ largest edge weight in $[\mathbf{W}_{iv},v\in\mathcal{N}_{i}]$. For each definition that requires a threshold, we first select a range of possible thresholds. For each threshold, we calculate the number of strong ties. Next, we compute the average number of strong ties across all thresholds. Finally, the threshold that produces the number of strong ties closest to this average is chosen as the final threshold. All the final thresholds can be found in the code. Based on these, we can generate the pseudo-ground-truth tie strength labels.
The comparisons among these definitions are shown in Table 1. As the semantic meaning of bidirectionality and weight information can be adjusted easily according to different datasets, we believe these 7 definitions can cover existing definitions of most prior works. Moreover, the evaluations from these definitions can reveal how the models will perform in real-world scenarios.
3.2 Comparison of Definitions
In this study, we collect three real-world social networks. We first introduce some of their details and then we investigate different tie strength definitions on these datasets. Specifically, we compare the strong/weak tie distribution, dive into the relationship between definitions through their correlations, and investigate the resilience of strong and weak ties according to diverse definitions in the Twitter dataset.
Curated Datasets
Table 2: The percentage of ties labeled as strong according to various tie strength definitions for each dataset.
Dataset $n$ $m$ I II III IV V VI VII
Bitcoin_alpha 3,682 12,976 0.75 0.21 0.17 0.11 0.45 0.26 0.23
CollgeMSG 1,899 13,838 0.47 0.24 0.21 0.11 0.36 0.21 0.17
Twitter 8,351 175,469 0.24 0.07 0.04 0.15 0.27 0.60 0.13
To cover different types of real-world social networks, we selected three datasets: Bitcoin_alpha (Derr et al. 2017) (a business transaction network), CollegeMSG (Panzarasa, Opsahl, and Carley 2009) (a college social network), and Twitter (Kheiri et al. 2023) (an online social network). These datasets range from public to private interactions and contain rich features for our experiments. Below are the dataset details:
Bitcoin_alpha: A cryptocurrency transaction network with anonymous online users as the nodes and their varying levels of trust/distrust ratings as directed weighted edges. We use text embeddings from user comments and filter out distrust edges to focus on trust-based strong/weak ties.
CollegeMSG: A social network of college students, where nodes represent users and edge weights are calculated based on the number of messages sent from one user to another.
Twitter: A following network where nodes are users and edges represent interactions (i.e., retweets and quotes) from one user to another. To ensure we are studying active users, we only include users that have at least posted or deleted a tweet in the prior 3 weeks. We also remove nodes that never interacted with other nodes to ensure the local threshold has practical meaning for them. The edge features are set to be the average embedding of the interacted tweet from source to target nodes.
In each of the resulting networks, we assigned an undirected tie between each pair of nodes with strength determined by the corresponding definitions. The tie weight features were obtained by averaging embeddings between two nodes for the Bitcoin_alpha and Twitter datasets, and by taking the sum of edge weights for CollegeMSG, as no semantic edge features are available for this dataset. This transformation prepares the dataset for the subsequent experiments.
Strong/Weak Tie Distribution
The definitions in Section 3.1 are applied to these datasets to create the pseudo ground-truth tie strength label. The ratios of pseudo ground-truth strong ties under each definition for each dataset are shown in Table 2. We can see that in most cases, these strong ties are the minority group in social networks. This is because strong ties usually serve as inter-community connections, and for most people, their active social circles are likely to be small in comparison with the whole network size. To further verify the effectiveness of our definitions, in Figure 1, we show the average weight, which is calculated as the sum of the weights of all edges between the endpoint nodes of the ties for both strong and weak ties. As shown in Figure 1, the average weight of strong ties is significantly higher than that of weak ties, demonstrating that our definitions successfully encapsulate the difference in strength between strong and weak ties.
Figure 1: Average weights of pseudo ground-truth of strong and weak ties. The average weight of strong ties is significantly higher than the average weight of weak ties.
(a) Bitcoin_alpha
(b) CollegeMSG
(c) Twitter
Figure 2: The Pearson correlation heatmaps (top row) and Euclidean distance heatmaps (bottom row) for Bitcoin_alpha, CollegeMSG, and Twitter under our definitions.
Correlation Between Definitions
We also use the heatmaps of Pearson correlation and Euclidean distance for these three datasets under various definitions to show their relationship. As shown in Figure 2, Definition I is distinct from other definitions as it is the only one that does not consider edge weight as a criterion. Definitions II and III are more similar to each other than to other definitions because they both rely on edge weight on a global scale. Additionally, they require a tie to meet a global threshold, which is set to make sure in most cases only a small fraction of ties can achieve. This makes such ties more likely to be bidirectional. This also stands for Definition VI and VII in our two interaction networks, Bitcoin_alpha and CollegeMSG datasets. However, the Twitter network is a following network based on a large online social network, where one direction followings are common (as shown in Definition I in Table 2). One user’s frequent interaction with another does not necessarily indicate mutual following (e.g. a huge fan of some celebrities, or active users who frequently comment on a topic), but this is much more rarely the case in Bitcoin_alpha and CollegMSG as people who don’t know another are less likely to trade or message frequently. This explains why Definition VI and VII are not so similar in the Twitter network as compared with the Bitcoin_alpha and CollegeMSG networks.
Strong/Weak Tie Resilience
Table 3: The ratio of pseudo ground-truth strong/weak ties in Twitter that became disconnected after $\Delta T$ weeks. R is the average of the ratio between weak and strong ties value.
$\Delta T=4$ $\Delta T=8$ $\Delta T=12$ R
Definition Strong Weak Strong Weak Strong Weak
I 0.82% 0.95% 1.37% 1.78% 1.92% 2.83% 1.31x
II 0.77% 0.93% 1.46% 1.69% 2.18% 2.64% 1.19x
III 0.79% 0.92% 1.38% 1.69% 1.99% 2.63% 1.24x
IV 0.83% 0.94% 1.54% 1.70% 2.36% 2.65% 1.12x
V 0.89% 0.93% 1.62% 1.70% 2.42% 2.67% 1.07x
VI 0.86% 1.01% 1.51% 1.93% 2.29% 3.09% 1.27x
VII 0.83% 0.93% 1.38% 1.72% 1.81% 2.73% 1.29x
R 1.14x 1.19x 1.30x 1.21x
Strong ties are expected to be less likely to break compared to weak ties. To demonstrate that our definitions align with this expectation, we use the Twitter dataset as an example and illustrate the disconnection rates of pseudo ground-truth strong and weak ties after $\Delta T$ weeks in Table 3. Overall, strong ties dissolved less often compared with weak ties across all definitions, and both strong and weak ties have more chances to dissolve with the increase of $\Delta T$. More specifically, Definitions I, III, and VII, which characterize strong ties as bidirectional, exhibit lower disconnection rates, affirming that bidirectional ties are more durable than unidirectional ones. Additionally, as $\Delta T$ increases, the probability of both strong and weak ties breaking rises. According to Figure 3, the ratio of weak to strong tie disconnection rates grows with $\Delta T$, indicating that weak ties break at a faster rate than strong ties. These findings not only corroborate our intuitive understanding but also offer valuable insights that (1) stronger ties are more resilient than weak ties, and (2) this tie resiliency ratio increases as the network evolves. These are important insights into tie strength and we believe that they can pave the way for future advancements in this field.
Figure 3: The disconnected rates ratio between weak over strong ties of each definition after $\Delta T$ weeks. Overall the ratio is becoming larger with the increase of $\Delta T$.
Table 4: Prediction accuracy of existing models on three datasets across seven definitions in the random weakly supervised setting. Overall models perform worse in Definition I, V, and VII, and MLP w/ $\mathbf{E}$ and GCN perform best.
Definitions Model Bitcoin_ahpha CollegeMSG Twitter Definitions Model Bitcoin_ahpha CollegeMSG Twitter
I (Avg. 0.720) Heuristic 0.695 0.572 0.739 V (Avg. 0.636) Heuristic 0.588 0.582 0.648
MLP w/o $\mathbf{E}$ 0.653 0.532 0.756 MLP w/o $\mathbf{E}$ 0.521 0.610 0.731
MLP w/ $\mathbf{E}$ 0.789 0.845 0.909 MLP w/ $\mathbf{E}$ 0.560 0.845 0.743
GCN 0.719 0.824 0.885 GCN 0.537 0.809 0.713
GTN 0.640 0.552 0.695 GTN 0.507 0.517 0.624
II (Avg. 0.852) Heuristic 0.749 0.737 0.910 VI (Avg. 0.713) Heuristic 0.679 0.751 0.827
MLP w/o $\mathbf{E}$ 0.794 0.758 0.930 MLP w/o $\mathbf{E}$ 0.741 0.781 0.573
MLP w/ $\mathbf{E}$ 0.783 1.000 0.939 MLP w/ $\mathbf{E}$ 0.726 0.900 0.544
GCN 0.794 0.976 0.931 GCN 0.674 0.863 0.637
GTN 0.789 0.757 0.931 GTN 0.650 0.794 0.558
III (Avg. 0.874) Heuristic 0.756 0.773 0.960 VII (Avg. 0.811) Heuristic 0.692 0.783 0.865
MLP w/o $\mathbf{E}$ 0.828 0.789 0.865 MLP w/o $\mathbf{E}$ 0.766 0.830 0.865
MLP w/ $\mathbf{E}$ 0.816 0.965 0.867 MLP w/ $\mathbf{E}$ 0.756 0.907 0.867
GCN 0.828 0.939 0.960 GCN 0.731 0.907 0.867
GTN 0.825 0.789 0.960 GTN 0.666 0.830 0.841
IV (Avg. 0.872) Heuristic 0.879 0.873 0.829 Avg Heuristic 0.720 0.724 0.825
MLP w/o $\mathbf{E}$ 0.889 0.882 0.843 MLP w/o $\mathbf{E}$ 0.742 0.740 0.808
MLP w/ $\mathbf{E}$ 0.882 0.890 0.845 MLP w/ $\mathbf{E}$ 0.759 0.907 0.830
GCN 0.889 0.905 0.845 GCN 0.739 0.889 0.834
GTN 0.886 0.894 0.845 GTN 0.709 0.733 0.779
4 Tie Strength Prediction Methods
In this section, we introduce the real-world settings and models that are used for later experiments.
4.1 Experiment Settings
Based on the definitions in Section 3.1, we can now assign the pseudo ground-truth tie strength in networks. However, as illustrated in Section 1, the ground-truth tie strength information is usually missing or incomplete in reality, and most online platforms lack mechanisms to distinguish between different types of ties. Therefore, when evaluating the existing tie strength prediction models’ performance, the common supervised learning setting with many labeled edges may not align well.
A more realistic setting would be weakly supervised settings, which stands for the ground-truth label is incomplete, or even unsupervised learning, where all ground-truth information is missing. While the traditional weakly supervised settings is to randomly select training edges from the overall network, a node-centric training selection might better align with the real-world social science domain; i.e., the idea of conducting surveys from a few individuals that provide their tie strengths (compared to randomly distributed user pairs).
Based on these thoughts, we proposed three experiment settings for this study:
  • Random weakly supervised setting: Randomly splitting the training edge set $\mathcal{E}^{\text{Tr}}$, validation edge set $\mathcal{E}^{\text{Val}}$, and testing edge set $\mathcal{E}^{\text{Test}}$, where $|\mathcal{E}^{\text{Tr}}|=|\mathcal{E}^{\text{Val}}|<<|\mathcal{E}^{\text{Test}}|$.
  • Node-centric weakly supervised setting: Randomly select a small number of nodes as hub nodes to form the training hub node set $\mathcal{V}^{\text{Tr}}_{\text{hub}}$ ($|\mathcal{V}^{\text{Tr}}_{\text{hub}}|<<|\mathcal{V}|$). Then randomly select the ties connected with these nodes to form the training edge set $\mathcal{E}^{\text{Tr}}$. In the remaining network, randomly select validation hub nodes set $\mathcal{V}^{\text{Val}}_{\text{hub}}$ such that $|\mathcal{V}^{\text{Val}}_{\text{hub}}|=|\mathcal{V}^{\text{Tr}}_{\text{hub}}|$, and get the validation edge set $\mathcal{E}^{\text{Val}}$ same as above. All the remaining edges in the network form the testing edge set $\mathcal{E}^{\text{Test}}$.
  • Unsupervised setting: In this setting, unsupervised models utilize the entire network without access to any pseudo ground-truth information regarding tie strength. The inferred tie strengths are then compared against the pseudo ground-truth labels to evaluate the models’ performance across various definitions.
Table 5: Prediction accuracy of existing models on three datasets across seven definitions in node-centric weakly supervised setting. Similarly, overall models perform worse in Definition I, V, and VII, and MLP w/ $\mathbf{E}$ and GCN perform best.
Definitions Model Bitcoin_ahpha CollegeMSG Twitter Definitions Model Bitcoin_ahpha CollegeMSG Twitter
I (Avg. 0.715) Heuristic 0.707 0.559 0.754 V (Avg. 0.631) Heuristic 0.509 0.563 0.597
MLP w/o $\mathbf{E}$ 0.746 0.506 0.663 MLP w/o $\mathbf{E}$ 0.526 0.638 0.735
MLP w/ $\mathbf{E}$ 0.833 0.817 0.914 MLP w/ $\mathbf{E}$ 0.533 0.813 0.736
GCN 0.741 0.817 0.757 GCN 0.516 0.798 0.733
GTN 0.650 0.511 0.743 GTN 0.495 0.588 0.685
II (Avg. 0.841) Heuristic 0.790 0.712 0.930 VI (Avg. 0.700) Heuristic 0.712 0.706 0.812
MLP w/o $\mathbf{E}$ 0.794 0.757 0.931 MLP w/o $\mathbf{E}$ 0.741 0.794 0.454
MLP w/ $\mathbf{E}$ 0.794 0.918 0.931 MLP w/ $\mathbf{E}$ 0.729 0.877 0.458
GCN 0.794 0.884 0.931 GCN 0.741 0.855 0.614
GTN 0.760 0.758 0.931 GTN 0.712 0.776 0.524
III (Avg. 0.873) Heuristic 0.821 0.751 0.960 VII (Avg. 0.812) Heuristic 0.724 0.745 0.867
MLP w/o $\mathbf{E}$ 0.828 0.789 0.960 MLP w/o $\mathbf{E}$ 0.766 0.829 0.867
MLP w/ $\mathbf{E}$ 0.828 0.969 0.960 MLP w/ $\mathbf{E}$ 0.766 0.883 0.867
GCN 0.828 0.885 0.960 GCN 0.766 0.832 0.853
GTN 0.805 0.789 0.960 GTN 0.736 0.830 0.853
IV (Avg. 0.867) Heuristic 0.861 0.881 0.799 Avg Heuristic 0.732 0.702 0.817
MLP w/o $\mathbf{E}$ 0.889 0.894 0.845 MLP w/o $\mathbf{E}$ 0.756 0.744 0.779
MLP w/ $\mathbf{E}$ 0.889 0.894 0.845 MLP w/ $\mathbf{E}$ 0.767 0.882 0.816
GCN 0.889 0.894 0.845 GCN 0.754 0.852 0.813
GTN 0.882 0.894 0.803 GTN 0.720 0.735 0.786
Note that for simplicity, the two weakly supervised settings are all transductive, and all edges in the training set are considered labeled. Inspired by the N-way K-shot setting in few-shot learning (Cheng et al. 2019), we explicitly set the size of the validation set to be equal (approximately equal in node-centric weakly supervised setting) to the size of the training set. To make sure the results between weakly supervised settings I and II are comparable, we explicitly set the number of training edges of these two settings to be approximately the same. For random weakly supervised setting, the training edge set sizes are 20 for Bitcoin_alpha, 70 for CollegeMSG, and 50 for Twitter. For node-centric weakly supervised setting, the number of hub nodes is set to 5 for Bitcoin_alpha, 4 for CollegeMSG, and 2 for Twitter, resulting in an average of 21, 68.3, and 52.3 training edges, respectively, across three splits.
4.2 Models for Tie Strength Prediction
As shown in Section 2.1, this study considers three main approaches to infer tie strength: heuristic methods, STC-based methods, and neural network methods. We focus on using structural features and edge features. For heuristic methods, we adopted the approach from  (Mattie et al. 2018) by training a Random Forest model (Breiman 2001) using structural features such as common neighbors, sum of end nodes’ degrees, sum of clustering coefficients, etc. For STC-based methods, we used STC-Greedy code from (Adriaens et al. 2018). For neural network methods, inspired by  (Zhuang et al. 2021), we selected two variants of a 3-layer Multi-layer Perception (MLP) (Taud and Mas 2018), one that includes edge features and one that does not, and two Graph Neural Networks (GNNs): GCN (Kipf and Welling 2017) and GTN (Shi et al. 2020), where one key difference is whether edge features are utilized in the message-passing process.
5 Empirical Study of Tie Strength Prediction
In this section, we present the results of existing models and offer an in-depth discussion. Following the setting of  (Gilbert and Karahalios 2009; Sohrabi and Akbari 2016), we adopted accuracy as the evaluation metric. Additionally, to reflect the intensity of strong and weak ties, similarly to (Sintos and Tsaparas 2014; Adriaens et al. 2018), we showed the average tie weight differences between strong and weak ties. All results are averaged over 3 runs with different data splits. The experiments were performed on a desktop with a 24GB NVIDIA GeForce RTX 3090 GPU and 64GB RAM. Note that we explicitly show the ”Avg” below each definition to reflect the average performance of the applied method, and use the last “Avg” section in Table 4 and  5 to reflect the average performance of each method on each dataset. Code and datasets are provided at https://github.com/XueqiC/tie˙strength˙prediction.
5.1 Random Weakly Supervised Setting
  • Nerual network models perform better: As shown in Table 4, the neural network methods outperform heuristic methods, with MLP w/ $\mathbf{E}$ achieving the best performance in the Bitcoin_alpha dataset, and GCN performing the best in CollegeMSG and Twitter datasets. In contrast, heuristic methods depict the worst overall performance in Bitcoin_alpha and CollegeMSG, and the second worst in Twitter. This highlights the greater overall effectiveness of neural network methods in leveraging structural and semantic features to predict tie strength, and heuristic methods that are built on pre-established rules may not be easily adaptable to different real-world scenarios. The fact that MLP w/ $\mathbf{E}$ depicts better than MLP w/o $\mathbf{E}$, but GCN performs better than GTN shows that leveraging semantic features has the potential to improve the performance, but GNNs need better mechanisms to integrate edge features.
  • Models struggle in the most distinct definitions: When analyzing the results across different definitions, we observe that the existing methods vary significantly in their performance. Specifically, Definitions I, V, and VI tend to have lower prediction accuracy. To analyze the reasons for these performance drops, we review the properties related to these definitions and find that these definitions are the most distinct cases according to Figure 2. More specifically, when looking into Definition I, we can find that performance is poor in Bitcoin_alpha and CollegeMSG but better in Twitter. This is likely because in Bitcoin_alpha and CollegeMSG, Definition I is less correlated with other definitions compared to Twitter. A similar pattern is seen in Definitions V and VI. This suggests that 1) there may be limitations in these definitions that cause them to assign pseudo ground-truth labels in a way that differs significantly from other definitions, and 2) model performance is highly dependent on the definition used, and methods struggle when the pseudo ground-truth labels are derived from a perspective different from the common understanding. When definitions reflect the most common understanding of what should be a strong/weak tie, models should be able to perform better.
  • Imbalance issue: In definitions with the highest overall performance, such as Definition III and IV, there are still concerns related to cases of extreme imbalance, like Twitter in Definition III and CollegeMSG in Definition IV. Based on the existing literature related to imbalance, the seemingly high prediction performance can be “unreal” as skewed distribution favors predictions toward the majority class, typically weak ties, which negatively impacts prediction accuracy for the minority class. We explore this imbalance issue further in Section 5.4.
  • Insiginificant average tie weight differences: The weight of a tie is the sum of all edge weights between its end nodes. Average weight differences between strong/weak ties across different datasets are shown in Table 6. As edge weight reflects the intensity of interactions or the level of trust, we expect the models can successfully reflect this by assigning strong labels to ties with relatively high edge weight. However, compared with the pseudo ground-truth, the existing methods fail to depict significant average weight differences across all definitions. Note that this implies that the existing methods fail to capture the assumed differences in strong versus weak ties having higher and lower edge weight, respectively. This could be because the models suffer from the imbalance issues, and/or are affected by noises in features and local structure information. More discussions are presented in Section 5.4.
5.2 Node-centric Weakly Supervised Setting
The prediction accuracy for node-centric random weakly supervised settings is presented in Table 5. Similar to the results in Section 5.1, neural network methods outperform heuristic methods in Bitcoin_alpha and CollegeMSG, with MLP w/ $\mathbf{E}$ achieving the best performance, while heuristic methods perform poorly in these datasets. However, heuristic methods show the best overall performance on Twitter, mainly due to their strong performance in Definition VI, where the predictors employed by heuristic methods successfully distinguish strong and weak ties. The poorer performance of neural network models here compared to Table 4 is likely because these models rely on hub nodes’ features and local class distribution information, which can introduce bias when this local information deviates significantly from global contexts.
Table 6: Random weakly supervised tie weight difference.
Method I II III IV V VI VII Avg.
Bitcoin_alpha
Ground Truth 1.658 6.215 6.227 5.872 2.966 3.960 3.985 4.698
Heuristic -0.04 0.791 0.521 0.958 -0.045 0.275 0.123 0.368
MLP w/o $\mathbf{E}$ 1.200 -3.489 -3.489 -3.489 -0.027 -3.489 -3.489 -2.467
MLP w/ $\mathbf{E}$ 0.212 -0.578 -0.578 -1.432 0.568 -1.856 -0.349 -0.573
GCN 1.107 -3.488 -3.488 -3.488 0.240 -1.156 -1.247 -1.788
GTN -0.011 0.314 0.458 0.148 0.089 -0.259 -0.337 0.058
Avg. 0.688 -0.872 -0.892 -0.905 0.399 -0.839 -0.902 -0.332
CollegeMSG
Ground Truth 5.323 10.253 10.392 14.048 6.699 10.217 11.132 9.866
Heuristic 0.461 1.737 2.014 0.576 -0.316 -1.511 -1.610 0.193
MLP w/o $\mathbf{E}$ 0.177 -4.311 -4.311 -3.022 -1.425 -3.018 -4.311 -2.889
MLP w/ $\mathbf{E}$ 5.360 10.268 10.449 10.279 9.594 12.322 13.166 10.777
GCN 6.009 10.191 11.764 2.634 9.403 7.061 13.052 8.302
GTN 1.664 -2.606 -4.320 -4.320 -0.953 -4.320 -4.320 -2.679
Avg. 3.832 3.755 4.665 3.699 3.667 3.392 4.185 3.885
Twitter
Ground Truth 0.458 2.499 3.143 0.532 0.641 0.100 0.635 1.287
Heuristic 0.125 0.411 0.739 0.078 0.064 -0.287 -0.064 0.152
MLP w/o $\mathbf{E}$0.071 -0.088 -0.101 -0.078 -0.036 0.002 -0.063 -0.062
MLP w/ $\mathbf{E}$ 0.124 0.478 -0.173 -0.173 0.642 0.086 -0.173 0.116
GCN 0.133 -0.173 -0.173 -0.173 0.144 -0.171 -0.173 -0.083
GTN -0.018 -0.173 -0.173 -0.173 -0.008 -0.054 -0.068 -0.095
Avg. 0.125 0.476 0.543 0.151 0.241 -0.054 0.016 0.214
In terms of performance across different definitions, we again observe lower accuracy in Definitions I, V, and VI, consistent with the findings in Section 5.1. This reinforces the observation that model performance depends on the definition used, and models struggle when the labels are derived from a perspective different from the others.
We also notice that models may suffer from imbalance issues, showing biased performance in cases of extreme imbalance, such as Twitter in Definition III and CollegeMSG in Definition IV. Moreover, the average tie weight differences across different datasets (as shown in Table 7) remain less pronounced compared to the pseudo ground-truth. We further discuss these in Section 5.4.
Table 7: Tie weight differences in the Node-centric setting.
Method I II III IV V VI VII Avg.
Bitcoin_alpha
Ground Truth 1.657 6.217 6.229 5.870 2.967 3.961 3.984 4.699
Heuristic -0.376 0.796 1.282 0.139 0.115 -0.101 -0.054 0.257
MLP w/o $\mathbf{E}$ 3.487 -3.487 -3.487 -3.487 0.009 -3.487 -3.487 -1.991
MLP w/ $\mathbf{E}$ 0.595 -3.489 -3.489 -3.489 0.312 0.211 -3.489 -1.263
GCN 2.153 -3.489 -3.489 -3.489 0.011 -3.489 -3.489 -1.897
GTN 1.056 0.186 0.261 -0.895 -0.057 -0.430 -0.489 -0.052
Avg. 1.429 -0.878 -0.948 -1.895 0.560 -0.889 -1.146 -0.481
CollegeMSG
Ground Truth 5.354 10.316 10.441 14.137 6.744 10.259 11.216 9.924
Heuristic 1.017 0.825 4.142 -0.405 -0.487 -0.163 0.786 0.816
MLP w/o $\mathbf{E}$ 0.182 -4.331 -4.331 -4.331 -4.331 -4.331 -4.331 -3.029
MLP w/ $\mathbf{E}$ 6.500 5.438 10.313 -4.331 8.458 11.558 7.218 6.165
GCN 6.166 6.087 5.236 -4.330 7.667 10.625 0.005 4.208
GTN -1.191 -0.780 -4.324 -4.324 -1.309 -3.226 -4.324 -2.783
Avg. 2.338 2.259 3.596 -1.764 2.957 3.453 1.381 2.460
Twitter
Ground Truth 0.458 2.499 3.143 0.531 0.641 0.100 0.635 1.287
Heuristic 0.544 0.815 0.253 0.081 0.089 -0.298 -0.106 0.197
MLP w/o $\mathbf{E}$0.049 -0.173 -0.173 -0.173 -0.173 -0.119 -0.173 -0.147
MLP w/ $\mathbf{E}$ 0.209 -0.173 -0.173 -0.173 0.839 -0.269 -0.173 0.012
GCN -0.173 -0.173 -0.173 -0.173 0.130 -0.029 -0.173 -0.109
GTN -0.050 -0.163 -0.163 -0.119 -0.033 0.050 -0.101 -0.085
Avg. 0.157 0.438 0.452 -0.004 0.249 -0.094 -0.015 0.169
(a) Datasets
(b) Definitions
Figure 4: Average weight differences of applying STC-Greedy compared with ground truth across different datasets and definitions. Overall, STC-Greedy depicts low performance and is not applicable for complex situations
5.3 Unsupervised Setting
The inference accuracy of using STC-Greedy is presented in Table 8. Since no label information is provided for training in this setting, the accuracy of STC-Greedy is understandably lower compared to the prediction accuracy discussed in the previous two sections. However, we still observe a notable performance drop in Definitions I, V, and VI, particularly in Bitcoin_alpha and CollegeMSG for Definitions I and V, and in Twitter for Definition VI. This aligns with our previous observations and further demonstrates that model performance is highly dependent on the definition used, with models struggling when pseudo ground-truth labels are derived from different perspectives.
The results for the average weight differences using STC-Greedy across different datasets (Figure 4(a)) and definitions (Figure 4(b)) show consistently insignificant differences compared to the ground truth. Note that since STC-Greedy cannot learn from the pseudo ground truth label from different definitions, their inference results are not changing. This also indicates that relying solely on STC properties cannot sufficiently highlight edges with larger weights, and better models are necessary to achieve improved performance.
Table 8: Accuracy of STC-Greedy in unsupervised setting.
Dataset I II III IV V VI VII Avg.
Bitcoin_alpha 0.295 0.742 0.769 0.819 0.540 0.697 0.718 0.654
CollegeMSG 0.529 0.721 0.751 0.844 0.616 0.756 0.787 0.715
Twitter 0.736 0.898 0.926 0.816 0.713 0.406 0.839 0.762
Avg. 0.520 0.787 0.815 0.826 0.623 0.620 0.781 0.710
Table 9: Apply quantity imbalance techniques in the random and node-centric weakly supervised setting.
CollegeMSG IV Twitter III
Methods Acc Macro-F1 Acc Macro-F1
Quantity reweight I
MLP w/o $\mathbf{E}$ 0.853 0.482 0.959 0.492
MLP w/ $\mathbf{E}$ 0.898 0.788 0.971 0.718
GCN 0.908 0.761 0.954 0.504
GTN 0.838 0.502 0.899 0.503
Quantity reweight II
MLP w/o $\mathbf{E}$ 0.878 0.479 0.935 0.492
MLP w/ $\mathbf{E}$ 0.820 0.709 0.967 0.640
GCN 0.859 0.749 0.960 0.491
GTN 0.893 0.473 0.915 0.493
Balanced hub nodes
MLP w/o $\mathbf{E}$ 0.870 0.490 0.953 0.493
MLP w/ $\mathbf{E}$ 0.908 0.798 0.964 0.581
GCN 0.887 0.777 0.959 0.492
GTN 0.825 0.503 0.949 0.496
5.4 Further Investigation
In this section, we conducted further experiments to discuss not only the concerns drawn from the experiments in Section 5.1 to Section 5.3 but also show the insights drawn from existing literature that is related to model performance.
Table 10: Average weight differences between strong and weak ties of MLP w/ $\mathbf{E}$ after applying quantity imbalance techniques is significantly improved.
CollegeMSG IV Twitter III
Ground Truth 14.048 3.143
Quantity reweight I 12.824 2.622
Quantity reweight II 9.587 4.142
Balanced Hub node 14.106 3.169
  • Imbalance Issue: As illustrated in Section 5.1 and  5.2, extreme imbalance cases exist such as CollegeMSG in Definition IV and Twitter in Definition III. To mitigate the imbalance issue, we adopt two quantity imbalance techniques to see if the performance can be improved.
    Quantity reweight (Huang et al. 2016) is one of the well-established quantity imbalance techniques, which assigns larger weight to minority classes during training. In this paper, for edge $e_{ij}\in\mathcal{E}^{Tr}$ that belongs to class $k$ (i.e., $\mathbf{Y}_{ijk}=1,k\in(S,W)$), the quantity reweight $w_{ij}^{q}$ that is applied to it during the training phase is:
    $w_{ij}^{q}=\frac{|\mathcal{E}^{\text{Tr}}|}{|\{\mathcal{E}_{k}^{\text{Tr}}\}_{k=1}^{C}|}$ (1)
    Besides quantity reweight, we also try to mitigate the issues by selecting hub nodes with balanced local class distribution in the node-centric weakly supervised setting.
    The results after applying the quantity imbalance techniques are presented in Table 9, where Quantity reweight I and II denote applying quantity reweight on weakly supervised settings I and II, respectively, and balanced hub nodes represent selecting hub nodes with balanced class distribution in the local neighborhood. We can find even though these quantity imbalance techniques can slightly improve the performance in some cases, i,e., MLP w/ $\mathbf{E}$ for CollegeMSG IV and Twitter III, they all fail to improve the performance accuracy in most other cases significantly. Besides, the Macro-F1 scores remain relatively low, indicating that the models still suffer from imbalance issues. These findings show that the current quantity imbalance techniques may not be sufficient for mitigating the imbalance issues in tie strength prediction tasks. As recent research in imbalanced node/graph classification (Chen et al. 2021; Song, Park, and Yang 2022; Wang et al. 2022) reveals a novel imbalance at the topology level and found its close correlation to the node/graph classification performance, we argue that there could also be a “topological issue” in tie strength prediction that stems from the uneven distributions of local class distributions around edges.
  • Insignificant Average Tie Weight Differences: The imbalance issue highlights that models struggle to differentiate between strong and weak ties. Strong ties typically have larger weights, and this imbalance can contribute to insignificant average weight differences. As demonstrated in Table 10, when applying the two imbalance techniques to MLP w/ $\mathbf{E}$, the average weight differences improve significantly. This suggests that using appropriate imbalance techniques can help to mitigate issues.
    However, these imbalance techniques rely on label information during training, making them unsuitable for unsupervised settings. An alternative improvement strategy is to develop models, such as graph contrastive learning, that can directly separate ties with smaller weights from those with larger weights without requiring label information.
    (a) STC-Greedy
    (b) MLP w/ $\mathbf{E}$
    (c) GCN
    Figure 5: Prediction accuracy of predicting tie strength in the Twitter dataset after $\Delta T$ weeks, based on the initial network.
    (a) Bitcoin_alpha
    (b) CollegeMSG
    (c) Twitter
    Figure 6: Average prediction accuracy of MLP w/ $\mathbf{E}$ and GCN in the random and node-centric weakly supervised setting across different training sample sizes and datasets across 10 runs.
  • Predicting Tie Strength in Dynamic Setting: Building on the work of  (Huang et al. 2018), which shows that tie strength can change over time, we propose that the observed tie strength today may differ from its strength in the future. This suggests that when we predict tie strength based on current data, we might actually be forecasting its future strength rather than its present state. To test this, we ran experiments using three methods (STC-Greedy, MLP w/ $\mathbf{E}$, and GCN) to predict tie strength in the initial Twitter network, comparing the predicted values with the pseudo ground-truth after $\Delta T$ weeks. As shown in Figure 5, the prediction accuracy changes over time for different prediction accuracy. We can observe that Definition VII increases as $\Delta T$ grows, suggesting that in this case, our predictions are more likely forecasting the future strength of a tie. This also implies that in this definition of friend and acquaintance, forming strong friendships takes time, even if conditions for a strong tie are present early on. In contrast, the prediction accuracy for Definition II and III decreases, suggesting that in those cases, predictions are more reflective of the current tie strength rather than its future development, and friendship in these definitions is more likely to be established very soon.
  • Influence of Experiment Setting: To get a deeper understanding of random and node-centric weakly supervised settings and how different models perform under these settings, we vary the size of training samples. The average accuracy of MLP w/ E and GCN on various datasets over 10 runs is shown in Figure 6. Overall, compared to their performance in node-centric supervised settings, models show improved performance in random weakly supervised settings. This might be because the node-centric setting may initially introduce biases based on selected hub nodes’ properties and connectivity patterns, potentially skewing the model’s understanding of the overall network structure. As training sizes increase, however, the gap in performance between the two settings decreases. This observation could indicate that the node-centric setting begins to include a broader variety of edge features and connection patterns. This reduces the initial bias, making the training set more representative of the network as a whole and aligning the performance of the node-centric approach with that of a random edge selection approach.
    Shifting the focus to model comparisons, performance differences between MLP w/ $\mathbf{E}$ and GCN vary across different datasets and training sizes. While GCN tends to outperform MLP w/ $\mathbf{E}$ in larger training scenarios, highlighting its ability to utilize local neighborhood patterns, this advantage is notably significant only in the CollegeMSG dataset. Therefore, further research is needed to understand the conditions under which MLP matches or exceeds GCN performance.
6 Discussion
Based on the previous sections, we conclude the performance of existing models and draw key insights from different perspectives in this section. The limitations and possible future directions are also proposed below:
  • Prediction Model Architectures: Overall, neural network models outperform heuristic methods. This is likely because neural network models are parametrized and can learn the complicated patterns in the data that are slipping from heuristic methods. When training sample size increases, neural network models that can effectively leverage the information in local neighborhoods are more likely to perform better in different scenarios. Besides, semantic features depict the potential to improve performance, but when incorporating them in GNNs, the aggregating and updating mechanism needs careful design. However, despite relatively good performance in weakly supervised settings, neural networks can not be directly implemented for unsupervised tasks as no training labels are provided.
    Also, we can conclude that even though heuristic models with predefined rules might outperform other baselines in biased training scenarios, they generally do not perform well in most other situations. This inconsistency demonstrates that the effectiveness of predefined rules is not guaranteed. For STC-Greedy, as it is not a data-driven learning method, it is hard to perform well in different complex scenarios. Moreover, none of the existing models can distinguish strong and weak ties to reflect the weight differences effectively.
    In this work, we seek to draw valuable insights based on the above observations that help guide the future directions in developing advanced tie strength prediction models. In the future, it would be worthwhile to focus on (1) developing better mechanisms to integrate semantic features into GNN models, (2) extending methods to enable neural networks to perform well even with biased training data by incorporating heuristic rules, and (3) designing graph contrastive learning models (You et al. 2020) that can help distinguish between large and small weight edges for weakly/unsupervised settings.
  • Tie Strength Definitions: Existing models are harder to perform well when pseudo ground-truth labels are derived from a perspective different from the common understanding of strong/weak ties, such as Definition I, V, and VI when the training sample size is relatively small. This also implies that there may be limitations in these definitions. Still, existing models are only good at learning to extract all-rounded information that reflects more of the definitions that align with our common sense. For future research, we can leverage this insight to select the definitions that reflect common understanding and should correspond to the most reasonable definition. Also, we can work on developing models that can perform better even for these distinct definitions, which will increase the models’ performance in different complex scenarios.
  • Weakly Supervised Experiment Setting: Overall, models tend to perform better in the random weakly supervised setting compared to node-centric (as expected), especially with less training data. Additionally, GNN models that effectively use local neighborhood information are more likely to perform well across different settings. One limitation of our study is that we did not thoroughly examine the reasons for performance differences across datasets, e.g., why MLP w/ $\mathbf{E}$ outperforms GCN in Bitcoin_alpha and Twitter but not in CollegeMSG with limited training samples. We believe these differences may stem from the distinct local neighborhood patterns in each dataset and further investigation could lead to models that perform better across a wider range of networks.
  • Tie Strength Imbalance Issues: In many cases, existing models suffer from imbalance issues that we observed cannot be consistently mitigated by quantity imbalance techniques. However, inspired by recent advances in imbalanced edge classification (Cheng et al. 2025), we believe that by examining the imbalance problem from local topological perspectives, an effective solution can be found and one future direction.
  • Predicting Tie Strength in Dynamic Networks: As noted in Section 5.4, different tie strength definitions allow us to infer not only the current tie strength but also its future state. Building on this observation and the insights from tie resilience in Section 3, one of the future research directions to strengthen our current finding could be focusing on developing methods to estimate tie resilience based solely on the current network structure.
7 Conclusions
In this study, we address the challenge posed by the absence of ground-truth tie strength labels and the varying interpretations of tie strength among researchers, which complicate the development of effective prediction methods for real-world applications. To tackle this, we conducted a comprehensive evaluation of existing tie strength prediction methods across a range of real-world definitions. We first categorized mainstream understandings of tie strength into seven standardized definitions and validated their relevance by analyzing class distributions and correlations. Through this analysis, we gained important insights into tie resilience from the tie dissolution perspective. We then conduct extensive experiments and analysis of the performance of existing methods and reveal key findings based on the models’ performance from model architecture and definitions perspectives. We also discovered that current methods suffer from imbalance issues that cannot be adequately addressed by existing quantity imbalance techniques, and different definitions of tie strength enable the prediction of tie strength in different stages of dynamic networks. Based on these, we propose strategies to enhance current models and more importantly, offer several promising directions for future research in this field.